OpenStudy (anonymous):

The sum of three consecutive EVEN integers is 42. What are the integers? (Name all three)

OpenStudy (anonymous):

42/3 = 14 So 14 + 14 + 14 = 42 14 - 2 = 12 14 + 2 = 16 Thus 12 + 14 + 16 = 42

OpenStudy (amoodarya):

name them as a,b,c so a=2k b=2(k+1)=2k+2 c=2(k+2)=2k+4 now a+b+c=42

OpenStudy (amoodarya):

2k + 2k+2 +2k+4=42 6k+6=42 --> k=6 so a=2*6 b=2*6+2=14 c=2*6+4=16

OpenStudy (anonymous):

another way to solve this is: x + x + 2 + x + 4 = 42 3x + 6 = 42 3x = 36 x = 12 12, 14, 16.